ID |
Date |
Author |
Group |
Subject |
Text |
 |
65
|
05 May 2025, 22:41 |
Tokushi Sato | SPB | Shift summary | X-ray
delivery
8.0 keV, 1300-1600 uJ pulse |
|
64
|
04 May 2025, 23:14 |
Tokushi Sato | SPB | Shift summary | X-ray
delivery
8.0 keV, 1300 uJ pulse |
|
63
|
03 May 2025, 23:13 |
Tokushi Sato | SPB | Shift summary | X-ray
delivery
8.0 keV, 1600 uJ pulse |
|
62
|
03 May 2025, 06:47 |
Tokushi Sato | SPB | Shift summary | **X-ray delivery**
- 8.3 , 1600 uJ pulse energy
**Optical laser delivery** |
|
61
|
02 May 2025, 08:07 |
Andreas Koch | XPD | Issue | Concerning eLog entry #59, SPB PopIn:
The problem of the motor position, |
|
60
|
01 May 2025, 23:35 |
Tokushi Sato | SPB | Shift summary | X-ray
delivery
8.3 / 9.3 keV, 1400 uJ |
|
59
|
01 May 2025, 07:25 |
Romain Letrun | SPB | Issue | SPB pop-in monitor was not dricing to the
switch when we retracted it to the "out
" position with the MDL this morning. |
|
58
|
30 Apr 2025, 22:46 |
Romain Letrun | SPB | Shift summary | X-ray delivery
9.3 keV, 1800 uJ pulse
energy
1 - 202 pulses at |
|
57
|
26 Apr 2025, 06:55 |
Tokushi Sato | SPB | Shift summary | Staff: PV, TS, JK + external collaborators
X-ray delivery:
10 keV, ~ 800 uJ, 128 pulses, tuning |
|
56
|
25 Apr 2025, 06:48 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
10 keV, ~ 850 uJ, 128 pulses, tuning
was needed for 1 hour |
|
55
|
24 Apr 2025, 08:42 |
Konopkova Zuzana | HED | Issue | self-seeding level in SA2 is currently
very low (~100 uJ, about 50% in the pedestal),
need more information on how to proceed - |
|
54
|
24 Apr 2025, 06:56 |
Jayanath Koliyadu | SPB | Shift summary | X-ray delivery:
10 keV, ~ 900 uJ, 128 pulses, tuning
was needed for few hours |
|
53
|
17 Apr 2025, 01:39 |
Chan Kim | SPB | Shift summary | X-ray delivery
18.8 keV, 260 uJ pulse
energy
Optical |
|
52
|
14 Apr 2025, 23:08 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
20 keV, 250-400 uJ, 250 pulses,
stable delivery |
|
51
|
14 Apr 2025, 09:50 |
Jan Gruenert | XPD | Status | Beam Downtime
During last night, apparently the
photon beam was down in all SASE beamlines |
|
50
|
13 Apr 2025, 22:55 |
Tokushi Sato | SPB | Shift summary | Staff: PV, TS, JK, FK, + users
X-ray delivery:
20 keV, ~ 300-400 uJ, 250 pulses, |
|
49
|
12 Apr 2025, 23:15 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
20 keV, ~ 500 uJ, 246 pulses, stable
delivery |
|
48
|
11 Apr 2025, 23:14 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
20 keV, ~ 400 uJ, 246 pulses,
stable delivery |
|
47
|
10 Apr 2025, 23:17 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
20 keV, ~ 400 uJ, 128 pulses, stable
delivery |
|
46
|
09 Apr 2025, 22:46 |
Tokushi Sato | SPB | Shift summary | X-ray delivery:
20 keV, ~ 400 uJ, 128 pulses, stable
delivery |
|
|